OpenStudy (anonymous):

Express the angle in radian measure. 108°= __π /__

OpenStudy (lochana):

\[180 = \Pi\] \[18 = \frac{Pi}{10}\] \[108 = \frac{6Pi}{10}\] \[108 = \frac{3Pi}{5}\]